OpenStudy (anonymous):

Calculus: Find the derivative. y=x/(square root of 2x-3)

OpenStudy (anonymous):

\[x \div \sqrt{2x-3}\]

OpenStudy (anonymous):

Do you know the divide rule for derivative ??

OpenStudy (anonymous):

low d high - high d low? :o

OpenStudy (anonymous):

It is : \[\frac{x}{\sqrt{2x - 3}}\]

OpenStudy (anonymous):

I would use product rule...

OpenStudy (anonymous):

Yes you can use whatever you want...

OpenStudy (anonymous):

can you show me how u get the answer because i did both rules and i didn't get the same answer from what my teacher gave me :\

OpenStudy (anonymous):

X*(2x-3)^(-1/2)

OpenStudy (anonymous):

Divide rule says : \[\frac{u}{v} = \frac{v \cdot u' - u \cdot v'}{v^2}\]

OpenStudy (anonymous):

Can you show your steps if you have solved from both the rules ??

OpenStudy (anonymous):

Uv'+vu' U=x U'=1 V= (2x-3)^(-1/2) V'= -(2x-3)^(-3/2)

OpenStudy (anonymous):

I erased it already xD

OpenStudy (anonymous):

Substitute, simplify .

OpenStudy (anonymous):

You can do it here.. I just want to see that how you have done it so that we can find the mistake..

OpenStudy (anonymous):

ok but product rule is better since there is only x at the top rite?

OpenStudy (anonymous):

Mine was product rule. Either works. Take your pick.

OpenStudy (anonymous):

Well you have negative exponents but it works fine.

OpenStudy (anonymous):

What is the difficulty you are facing in using divide rule ??

OpenStudy (anonymous):